how do you find the value of x?
its a dumb question
by i still dont get the right answer
my cousin gave me the answer
but i dont know how to do it can u
explain it step by step.lol
its..2(2x+12)+6x-20=180
my answer is 23..
am i right?

Answer by stanbon(75887)

2(2x+12)+6x-20=180
Divide thru by 2 to get:
2x+12 + 3x-10=90
5x+2=90
5x=88
x=17.6
Cheers,
Stan H.